If the growth rate of the number of television in a country is 1% and the growth rate of the country's population is 3%, the growth rate of the number of television per-capita is:

A)2%

b)-2%

c)4%

d)-4%
Answer #1

At base assume both population and television is 100 units. Given the growth rates of 1% and 3% for television and population respectively, the television and population will be 101 and 103 respectively and so own

Thus In year 0, television = 100,    population = 100

In year 1,    television = 100*1% = 101, population = 100*3% = 103

In year 2, television = 101*1% =102.01 population = 103*3% =106.09

Television par-capita in year 1 = 101/103 = 0.9805

Television per-capita in year 2= 102.01/106.09 = 0.961

Growth rate of per capita = [(0.961- 0.9805)/0.9805] = -0.019*100/100 = 2% ( approx).